OpenStudy (anonymous):

Write an equation of the sine function. amplitude: 4, period: pi, phase shift: -1/6 pi, vertical shift: -1

OpenStudy (anonymous):

@ejune420 What do you mean by amplitude here? sin can take maximum value of +1

OpenStudy (kropot72):

The general equation is as follows: \[y=A \sin (Bx+C)+D\] where: amplitude is |A| period T = 2 pi/B phase shift = C/B 'y' shift = D In your question the term A in the general equation can be either 4 or -4. \[T=\pi=\frac{2 \pi}{B}\ .........(1)\] You need to solve equation (1) to find B. \[phase\ shift=-\frac{1}{6} \pi=\frac{C}{B}\ .........(2)\] You need to substitute the value for B into equation (2) and then solve to find C. \[D=-1\]